pH of diluted orange juice

Reading on the Burette (mL)

4.15

0.50

4.19

1.00

4.19

1.50

4.22

2.00

4.31

2.50

4.35

3.00

4.76

3.50

4.87

4.00

4.97

4.50

5.09

5.00

5.23

5.50

5.37

6.00

5.52

6.50

5.69

7.00

5.83

7.50

5.98

8.00

6.06

8.25

6.11

8.50

6.22

8.75

6.32

9.00

6.43

9.25

6.54

9.50

6.68

9.75

6.89

10.00

7.08

10.25

7.40

10.50

7.87

10.75

8.24

11.00

8.63

11.25

10 mL of orange juice diluted with 75 mL of distilled water and titrated with 0.1M NaOH and 3 drops of phenolphtalein. Calculate % acid of the juice.
Answer #1

The data is plotted:

The first equivalence point occurs at 3.25 mL, the concentration of citric acid in orange juice is calculated:

Ca = Cb * Vb / Va = 0.1 M * 3.25 mL / 10 mL = 0.0325 M

The percentage of acid in solution is calculated:

% acid = 0.0325 mol / L * (192.12 g / 1 mol) * (1 L / 1000 mL) * 100 = 0.62%

If you liked the answer, please rate it positively, you would help me a lot, thanks.